OpenStudy (anonymous):

Solve the equation: 5x(x + 3) = 0

OpenStudy (anonymous):

1 plus 1 equals 2

OpenStudy (anonymous):

? please dont post stupid stuff unless it is going to help me. i dont have time for it.

OpenStudy (anonymous):

its educational

OpenStudy (anonymous):

btw Expand the brackets: 5x(x+3) = 0 5x^2+15x=0 Then subtract both sides by 15x 5x^2 = -15x Dive both sides by 5x x = -3 ( You see the -3 would cancel out the other 3 inside the bracket to give you 0. (x could also be 0 to make 5x(x+3) = 0 ) now call me stupid...smh

OpenStudy (anonymous):

Yea i got it already thaks. i didnt call u stupid. B/

OpenStudy (anonymous):

alright ;)

OpenStudy (anonymous):

Ok, Do you know this one? It would be highly appreiated :D Solve the equation for x: 3x2 = 6x

OpenStudy (anonymous):

isolate x... 3x^2-6x = 0 you can use factoring here.. 3x(x-2) = 0 3x = 0 x - 2 =0 so x = 0 x=2